# Average-Case Deterministic Query Complexity of Boolean Functions with Fixed Weight

### v2

@@ 5 Circuits
$\geq$ $\ n\left(1-\frac{O(\log n)}{w}\right)(1-ph)$ $\Big{(}1-ph\leq(1-p)^{h}\Big{)}$
$=$ $\ n\left(1-\frac{\log n}{\Omega(w)}\right).$ $\left(ph\leq\frac{3}{w}\left(1+\frac{w}{n}\right)\right)$ (15)
On the other hand, we can query $f$ by evaluating all the clauses one by one. By Lemma 4.1, $\mathrm{D}_{\mathrm{ave}}(g)\leq\log m+2\leq w-\log n+4$. Thus,
$\mathrm{D}_{\mathrm{ave}}(f)\leq h\cdot\mathrm{D}_{\mathrm{ave}}(g)\leq n\left(1-\frac{\log n-4}{w}\right)=n\left(1-\frac{\log n}{O(w)}\right).$ (16)
@@ 6 Conclusion
In this paper, we study the average-case query complexity of boolean functions under the uniform distribution. We give an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$ in terms of its weight. We prove, for almost all fixed-weight boolean functions, our upper bound is tight up to an additive logarithmic term). We show that, for any $w\geq\log n+\sqrt{\log n}$, there exists a DNF formula of width $w$ and size $2^{w}/w$ such that $\mathrm{D}_{\mathrm{ave}}(f)=n(1-\frac{\log n}{\Theta(w)})$, which suggests that the switching lemmas are almost tight. We also study or compute $\mathrm{D}_{\mathrm{ave}}(f)$ of several functions, including the penalty shoot-out function, symmetric functions, linear threshold functions and the tribes functions.
Theorems 1.1 and 1.2 essentially relate $\mathrm{D}_{\mathrm{ave}}(f)$ to the zero-order Fourier coefficient $\widehat{f}(\{\emptyset\})$. Establishing an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$ in terms of higher-order Fourier coefficients (e.g. influences) can deepen our understanding of $\mathrm{D}_{\mathrm{ave}}(f)$ and pave the way for more applications. For example, it is unknown whether the lower bound $\mathrm{D}_{\mathrm{ave}}(\mathrm{Hex}_{L\times L})\geq L^{1.5+o(1)}$, which coincides with the expected optimal playing time for the random-turn Hex game, is tight or not [PSS+07], and such a bound on $\mathrm{D}_{\mathrm{ave}}(f)$ might shed light on the question.
It remains open to prove the tightness of the upper bounds (Corollary 5.7, 5.8) for CNF/DNF formulas. If true, it would imply that the switching lemmas are tight.
@@ Acknowledgements
Appendix A General lower bounds
Theorem A.1 (OSSS inequality [OSS+05]).
For any decision tree $T$ computing a boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$,
$\sum_{i=1}^{n}\delta_{i}\cdot\mathrm{{Inf}}_{i}[f]\geq\mathrm{{Var}}[f],$
where $\delta_{i}\in[0,1]$ is the probability that the $i$-th variable $x_{i}$ is queried by the decision tree $T$. In particular, taking $T$ as an optimal tree computing $f$ in terms of $\mathrm{D}_{\mathrm{ave}}(f)$, we have
$\mathrm{D}_{\mathrm{ave}}(f)=\sum_{i=1}^{n}\delta_{i}\geq\frac{\mathrm{{Var}}[f]}{\max_{i\in[n]}\mathrm{{Inf}}_{i}[f]}.$
Theorem A.2 (OS inequality [OS06]).
For every boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$,
$\left(\sum_{i=1}^{n}\widehat{f}(\{i\})\right)^{2}\leq\mathrm{D}_{\mathrm{ave}}(f)\leq\log\mathrm{DT}_{\mathrm{size}}(f),$
where $\mathrm{DT}_{\mathrm{size}}(f)$ is the decision tree size. In particular, $\mathrm{D}_{\mathrm{ave}}(f)\geq\mathrm{s}_{\mathrm{ave}}(f)^{2}$ when $f$ is monotone.
Appendix B Special function classes
B.1 AND, OR and XOR
Define $\mathrm{AND}_{n}(x)=\bigwedge_{i=1}^{n}x_{i}$ and $\mathrm{OR}_{n}(x)=\bigvee_{i=1}^{n}x_{i}$ and $\mathrm{XOR}_{n}(x)=\bigoplus_{i=1}^{n}x_{i}$. The following proposition gives the exact value of $\mathrm{D}_{\mathrm{ave}}(f)$ for boolean functions $f$ with weight $1$, including the AND function. It shows a super-exponential gap between $\mathrm{D}(f)$ and $\mathrm{D}_{\mathrm{ave}}(f)$, since $\mathrm{D}(\mathrm{AND}_{n})=n$ and $\mathrm{D}_{\mathrm{ave}}(\mathrm{AND}_{n})=O(1)$, and also gives the exact value of $\mathrm{D}_{\mathrm{ave}}(\mathrm{OR}_{n})$.
Proposition B.1.
Proof.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function with a unique black point $z\in\{0,1\}^{n}$. We prove $\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})$ by induction on $n$. When $n=1$, it is easy to verify $\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})=1$.
$=1+\frac{1}{2}\cdot 2\left(1-\frac{1}{2^{n-1}}\right)$
$=2\left(1-\frac{1}{2^{n}}\right),$
where the second step is by the induction hypothesis. ∎
Corollary B.2 (Exercise 8.65 in [ODo14]).
$\mathrm{D}_{\mathrm{ave}}(\mathrm{AND}_{n})=\mathrm{D}_{\mathrm{ave}}(\mathrm{OR}_{n})=2(1-\frac{1}{2^{n}})$.
Proof.
By Proposition 3.6, $\mathrm{D}_{\mathrm{ave}}(\mathrm{AND}_{n})=\mathrm{D}_{\mathrm{ave}}(\mathrm{OR}_{n})$, since the OR function is the dual of the AND function. Then, the corollary follows from Proposition B.1. ∎
In addition, we show the XOR function (and its negation) is the only function with the maximum complexity $\mathrm{D}_{\mathrm{ave}}(f)=n$.
Proposition B.3.
Let $f$ be an $n$-variable boolean function. Then, $\mathrm{D}_{\mathrm{ave}}(f)=n$ if and only if $f=\mathrm{XOR}_{n}$ or $f=\mathrm{XOR}_{n}\oplus 1$.
Proof.
If $f$ is $\mathrm{XOR}_{n}$ or $\mathrm{XOR}_{n}\oplus 1$, then $\mathrm{D}_{\mathrm{ave}}(f)=n$, since $f|_{\rho}$ is not constant for any restriction $\rho$ fixing $\leq n-1$ variables.
Now we use induction to prove the opposite direction that if $\mathrm{D}_{\mathrm{ave}}(f)=n$, then $f$ must be $\mathrm{XOR}_{n}$ or $\mathrm{XOR}_{n}\oplus 1$. When $n=1$, the conclusion is obvious. Assuming the conclusion is true for $n-1$, we prove it for $n$. Suppose an optimal query algorithm computing $f$ is to query the variable $x_{i}$ first. Since $\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=b})\leq n-1$ for any $b\in\{0,1\}$, we have
$n=\mathrm{D}_{\mathrm{ave}}(f)\leq 1+\frac{1}{2}\left(\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=0})+\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=1})\right)\leq n,$
which implies that $\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=0})=\mathrm{D}_{\mathrm{ave}}(f|_{x_{i}=1})=n-1$. By the induction hypothesis, for any $b\in\{0,1\}$, we know $f|_{x_{i}=b}$ is either $\mathrm{XOR}_{n-1}\text{ or }\mathrm{XOR}_{n-1}\oplus 1$.
If $f|_{x_{i}=0}=f|_{x_{i}=1}$, meaning $f=f|_{x_{i}=0}$, then the variable $x_{i}$ is not pivotal on any input, which contradicts our assumption that $\mathrm{D}_{\mathrm{ave}}(f)=n$. Thus, $f|_{x_{i}=0}$ and $f|_{x_{i}=1}$ are not the same. We conclude that $f=\mathrm{XOR}_{n}$ if $f|_{x_{i}=0}=\mathrm{XOR}_{n-1}$, and that $f=\mathrm{XOR}_{n}\oplus 1$ if $f|_{x_{i}=0}=\mathrm{XOR}_{n-1}\oplus 1$. ∎
B.2 Penalty shoot-out function
@@
In the worst case, the winner cannot be declared until the last round is finished. So $\mathrm{D}(\mathrm{PSO}_{n})=2n+1$.
In the average case, we prove $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=4-\frac{3}{2^{n}}$ by induction on $n$. When $n=0$, $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=1$, because one kicked by team A decides the game. Assuming $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n-1})=4-\frac{3}{2^{n-1}}$ holds, let us prove $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=4-\frac{3}{2^{n}}$. In each round, there are four cases:

### v3

@@ 4 DNFs, circuits, and formulas
$\geq$ $\ n\left(1-\frac{\log n+O(\log w)}{w}\right)(1-ph)$ $\Big{(}1-ph\leq(1-p)^{h}\Big{)}$
$=$ $\ n\left(1-\frac{\log n}{\Omega(w)}\right).$ $\left(ph\leq\frac{1}{w}\right)$ (16)
On the other hand, we can query $f$ by evaluating all the clauses one by one. By Lemma 3.2, $\mathrm{D}_{\mathrm{ave}}(g)\leq\log m+2\leq w-\log n+2$. Thus,
$\mathrm{D}_{\mathrm{ave}}(f)\leq h\cdot\mathrm{D}_{\mathrm{ave}}(g)\leq n\left(1-\frac{\log n-2}{w}\right)=n\left(1-\frac{\log n}{O(w)}\right).$ (17)
@@ 5 Conclusion
In this paper, we studied the average-case query complexity of boolean functions under the uniform distribution. We prove an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$ in terms of its weight; on the other hand, we prove that for almost all fixed-weight boolean functions, the upper bound is tight up to an additive logarithmic term. We show that, for any $w\geq 2\log n$, there exists a DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$ such that $\mathrm{D}_{\mathrm{ave}}(f)=n(1-\frac{\log n}{\Theta(w)})$, which suggests that the criticality bounds $O(w)$ and $O(\log s)$ are tight up to a multiplicative $\log n$ factor.
Theorems 1.1 and 1.2 essentially relate $\mathrm{D}_{\mathrm{ave}}(f)$ to the zero-order Fourier coefficient $\widehat{f}(\{\emptyset\})$. Establishing an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$ in terms of higher-order Fourier coefficients (such as influences) would be valuable. For example, it is unclear whether the lower bound $\mathrm{D}_{\mathrm{ave}}(\mathrm{Hex}_{L\times L})\geq L^{1.5+o(1)}$ is tight [PSS+07]; bounding $\mathrm{D}_{\mathrm{ave}}(f)$ in terms of Fourier coefficients might shed light on the open problem.
It remains open to prove tight upper bounds on $\mathrm{D}_{\mathrm{ave}}(f)$ for $k$-DNF, as well as for bounded depth formulas and circuits.
@@ Acknowledgements
Appendix A Penalty shoot-out function
@@
In the worst case, the winner cannot be declared until the last round is finished, so $\mathrm{D}(\mathrm{PSO}_{n})=2n+1$.
We prove $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=4-\frac{3}{2^{n}}$ by induction on $n$. When $n=0$, $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=1$, because one kick by team A decides the game. Assuming $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n-1})=4-\frac{3}{2^{n-1}}$ holds, let us prove $\mathrm{D}_{\mathrm{ave}}(\mathrm{PSO}_{n})=4-\frac{3}{2^{n}}$. In each round, there are four cases: